OpenStudy (anonymous):

Simplify: log (1/2) 8

OpenStudy (anonymous):

(1/2)^x = 8 x = -3 log(1/2) 8 = -3

OpenStudy (anonymous):

does "simplify" mean "evaluate?

OpenStudy (anonymous):

Yes I think so

OpenStudy (anonymous):

evaluate means find a value

OpenStudy (anonymous):

Then no, it's means to simplify it as far as it can go

OpenStudy (anonymous):

if so your job is to answer this question; one half to what power is 8, or "what would i raise one half to , to get the number 8?"

OpenStudy (anonymous):

in other word solve for y \[(\frac{1}{2})^y=8\]

OpenStudy (anonymous):

Then no, it's means to simplify it as far as it can go

OpenStudy (anonymous):

Ohkay, how do I do that??

OpenStudy (anonymous):

and the answer is -3 pretty much by inspection

OpenStudy (anonymous):

2 cubed is 8, and you need to flip it

OpenStudy (anonymous):

you have to take the reciprocal and then cube. you pretty much have to figure it out from scratch, there is no formula.

OpenStudy (anonymous):

but if you have a calculator i can show you how to cheat

OpenStudy (anonymous):

Then no, it's means to simplify it as far as it can go

OpenStudy (anonymous):

Ohh I see. That makes sense

OpenStudy (anonymous):

it is easy to do if you are allowed to use a calculator

OpenStudy (anonymous):

You should show me how to cheat :) lol

OpenStudy (anonymous):

ok your calculator already has a log on it. it just says "log"

jimthompson5910 (jim_thompson5910):

\[\large \log_{\frac{1}{2}}\left(8\right)\] \[\large \frac{\log_{10}\left(8\right)}{\log_{10}\left(\frac{1}{2}\right)}\] \[\large \frac{\log_{10}\left(2^{3}\right)}{\log_{10}\left(2^{-1}\right)}\] \[\large \frac{3\log_{10}\left(2\right)}{-1\log_{10}\left(2\right)}\] \[\large \frac{3}{-1}\] \[\large -3\] So \[\large \log_{\frac{1}{2}}\left(8\right)=-3\]

OpenStudy (anonymous):

yeah i know how to use log i just don't know how to plug in th sub (1/2) thing

OpenStudy (anonymous):

type in \[\log(8)\div\log(\frac{1}{2})\] and you will get your answer

OpenStudy (anonymous):

You cannot. You have to change the base

OpenStudy (anonymous):

Base! That was the word I was looking for. Does that trick work with every problem or just ones that it happens to work on like this one?

OpenStudy (anonymous):

you can find the log of any number for any base using this \[\log_b(x)=\log(x)\div\log(b)\] where log means the log on your calculator

OpenStudy (anonymous):

Then no, it's means to simplify it as far as it can go

OpenStudy (anonymous):

Okay cool. I think I remember that formula

OpenStudy (anonymous):

if i want \[\log_7(95)\] i just type \[\log(95)\div\log(7)\] and i am done

OpenStudy (anonymous):

\[log_b (a) = k \implies b^k = a\] Taking the log (base 10) of both sides we have: \[log(b^k) = log(a)\]\[\implies k\cdot log(b) = log(a)\]\[\implies k = \frac{log(a)}{log(b)}\] But we started with the premise that \(log_b(a) = k\). Therefore: \[log_b(a) = \frac{log(a)}{log(b)}\]

OpenStudy (anonymous):

so rather than spending all my time figuring out this to the that is the other, i can just use the "change of base" formula, which is what polpak wrote

OpenStudy (anonymous):

Yep. Though I cannot usually remember the formula, so I have to derive it every time ;p

OpenStudy (anonymous):

i don't believe you. you want to solve \[b^x=A \]for x i be you go right to \[x=\frac{\log(A)}{\log(b)}\] without thinking

OpenStudy (anonymous):

Nope. I'd have to plod along (I'd probably use ln too) \[\implies ln(b^x) = ln(A) \implies x = \frac{ln(A)}{ln(B)}\]

OpenStudy (anonymous):

err B=b But even then this isn't the situation we're looking at. We're talking about \[x = log_b45\]

OpenStudy (anonymous):

And yeah, I'm gonna have to go to the derivation for that also

OpenStudy (anonymous):

Wonderful, I wrote that down in my notes so I don't forget.